# Global ${\mathcal{L}}^{2}$ minimization at uniform exponential rate via geometrically adapted gradient descent in Deep Learning

### v4

@@ 4.2. Perturbation away from local equilibrium
$\underline{\theta}_{0}^{\prime}=\underline{\theta}_{*}+\kappa V_{\underline{\theta}_{*}}$ (4.15)
@@
$V_{\underline{\theta}_{*}}=\nabla_{\underline{\theta}}g_{\alpha}\big{|}_{\underline{\theta}=\underline{\theta}_{*}}$ (4.16)
with $g_{\alpha}$ defined as in (4.13), and a suitable choice of $\alpha\in\{1,\dots,r\}$. The sign of the parameter $\kappa\in{\mathbb{R}}$ is chosen such that the cost is decreasing,
@@ 6. The borderline case $K=QN$
Acknowledgments: The author thanks Patricia Muñoz Ewald for discussions, and gratefully acknowledges support by the NSF through the grant DMS-2009800, and the RTG Grant DMS-1840314 - Analysis of PDE.

### v5

@@ 4.2. Perturbation away from local equilibrium
$\underline{\theta}_{0}^{\prime}=\underline{\theta}_{*}+V_{\underline{\theta}_{*}}$ (4.15)
@@
$V_{\underline{\theta}_{*}}=\kappa_{\alpha}\nabla_{\underline{\theta}}g_{\alpha}\big{|}_{\underline{\theta}=\underline{\theta}_{*}}$ (4.16)
with $g_{\alpha}$ defined as in (4.13), and a suitable choice of $\alpha\in\{1,\dots,r\}$. The sign of the parameter $\kappa_{\alpha}\in{\mathbb{R}}$ is chosen such that the cost is decreasing,
@@
More generally, one may choose a linear superposition of vectors of the form (4.16), for each of which one obtains (4.17).
@@ 6. The borderline case $K=QN$
Acknowledgments: The author thanks Patricia Muñoz Ewald for discussions, the anonymous referees for helpful comments, and gratefully acknowledges support by the NSF through the grant DMS-2009800, and the RTG Grant DMS-1840314 - Analysis of PDE.
There are no financial or non-financial competing interests that are directly or indirectly related to this work.